Q2) A S 1,000 bond with an 8% coupon rate, with coupons paid semi-annually, is maturing in 10 years. If the quoted YTM is 10%, what is the bond price?
Answer #1

Number of periods = n = 10*2 = 20 semiannual periods

Yield I/Y= 10%/2 = 5%

Semiannual Payment P = 8%*1000/2 = $40

Face Value FV = $1000

Hence, PV = P/(1+r) + P/(1+r)2 + .... + P/(1+r)n + FV/(1+r)n

= P[1 - (1+r)-n]/r + FV/(1+r)n = 40(1 - 1.05-20)/0.05 + 1000/1.0520 = $875.38

If we use the financial calculator, we get PV = $875.38